OpenStudy (micahm):

i don't know if i have ask this question already The perimeter of a rectangular flower garden is 60 m and its area is 225 m2. Find the length of the garden how do i show my work and solve this can i have step by step help

OpenStudy (solomonzelman):

\(\huge\color{magenta}{ \rm L }\) \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \color{lightcyan}{\Huge\frac{~~~~~~~~\frac{~}{~~~\frac{~\frac{}{}~}{~\frac{~}{~}}~~~~}~~~}{~\frac{\frac{~}{~\frac{~\frac{}{~}}{~}}~~}{~}~~} } }}}\) \(\huge\color{darkviolet}{ \rm W }\) \(\huge\color{blue}{ \rm A_{rectangle} = \color{magenta}{ \rm L }\times\color{darkviolet}{ \rm W } }\) is the formula/concept: ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ In your case, \(\huge\color{magenta}{ \rm L }\) \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \color{lightcyan}{\Huge\frac{~~~~~~~~\frac{~}{~~~\frac{~\frac{}{}~}{~\frac{~}{~}}~~~~}~~~}{~\frac{\frac{~}{~\frac{~\frac{}{~}}{~}}~~}{~}~~} } }}}\) \(\huge\color{darkviolet}{ \rm 60 }\) \(\huge\color{blue}{ \rm 225 = \color{magenta}{ \rm L }\times\color{darkviolet}{ \rm 60 } }\)

OpenStudy (solomonzelman):

can you solve for \(\LARGE\color{magenta}{ \rm L }\) ?

OpenStudy (micahm):

ok so do i mulp

OpenStudy (igreen):

\(\sf 225 = L \times 60 \rightarrow 225 = 60L\)

OpenStudy (mathmath333):

cool drawing :)

OpenStudy (igreen):

mulp..?

OpenStudy (igreen):

Just solve for 'L': \(\sf 225 = 60L\)

OpenStudy (micahm):

13500

OpenStudy (solomonzelman):

thanx math:) just a little trick with a box

OpenStudy (igreen):

No..how did you get that..?

OpenStudy (micahm):

i mulp 225x60

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \rm 225=60 \times \color{magenta}{L} }\) isolate the \(\LARGE\color{black}{ \color{magenta}{L} }\)

OpenStudy (igreen):

@SolomonZelman Yes, me likes your rectangle :D

OpenStudy (solomonzelman):

lol, ty!

OpenStudy (igreen):

No, we do not multiply.. To get 'L' by itself, we must get rid of the '60' that is being multiplied to it..and to do that we do the opposite..so what's the opposite of multiplication? @micahm

OpenStudy (micahm):

divd

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

you divide:)

OpenStudy (micahm):

3.75

OpenStudy (igreen):

Yep, you got it.

OpenStudy (solomonzelman):

yes, that is correct.

OpenStudy (solomonzelman):

didn't have a cac, so entered in wolfram to check

OpenStudy (igreen):

Google works too.. :P @SolomonZelman

OpenStudy (solomonzelman):

never think of google, dky

OpenStudy (micahm):

so 3.75 is the answer to the problem

OpenStudy (solomonzelman):

yes, but your unit, is meter, don't foget

OpenStudy (solomonzelman):

forget

OpenStudy (igreen):

Lol, all you have to do is plug in '225 / 60' in google where you usually have your URL..it pops right up..you don't even have to press 'Enter'.

OpenStudy (solomonzelman):

I like that they are using an SI unit in the example, lol :D

OpenStudy (solomonzelman):

alrighty, yw:)

OpenStudy (igreen):

yep..3.75 is your answer..

OpenStudy (micahm):

thank you very much both of you

OpenStudy (solomonzelman):

:D